# DAViD: Modeling Dynamic Affordance of 3D Objects using Pre-trained Video Diffusion Models

### v2

@@ 4.3 Qualitative Results
4D HOI Generation for Various Objects. Different from previous studies, we leverage the pre-trained video diffusion model, which allows to learn dynamic patterns of HOI for unbounded object categories objects. Fig. 4 visualizes the HOI motion for various objects generated by DAViD. The frames are visualized in temporal order. The results show that our method effectively models the dynamic patterns in HOIs. In Fig. 4 (a), Fig. 4 (b), Fig. 4 (c) we observe that the human motion and the relative position of the object are well modeled in relatively static scenarios. In cases with more dynamic human motion, we find that various interaction patterns, including pulling--Fig. 4 (d), pushing--Fig. 4 (h), lifting--Fig. 4 (g), and holding--Fig. 4 (f) are well aligned with the typical use of objects of humans. Also, we can observe that our method can effectively capture complex human motions and their corresponding object motions, such as back and forth movement, as shown in Fig. 4 (e) and Fig. 4 (j). Importantly, unlike previous studies where motion translation typically resulted from walking, we observe that our model effectively captures cases of translation changes without foot movement, as shown in Fig. 4 (i).
Figure 5: Combining Concepts. Our HOI motion concepts can be combined with original knowledge of pre-trained MDM, even for the multiple HOI motion concepts.
@@
[…]
@@
[…]
@@ 4.4 Quantitative Results
Comparison with Baselines. To verify the efficacy of our model, we compare DAViD against OMOMO [50] and CHOIS [51] on the FullBodyManipulation Dataset [50]. Specifically, we measure the performance of DAViD on each of the test datasets of the FullBodyManipulation [50], selected from baselines. For the metrics about HOI quality, hand contact is considered when the average distance between the 25 hand joints of SMPL-X [69] and the object vertices is smaller than a threshold, 0.06. As shown in Tab. 1, DAViD outperforms the baselines for all contact metrics. This shows that DAViD generates plausible HOI motion, demonstrating the efficacy of our approach in modeling the object pose distribution from the human pose, and sampling object motion using TGS. Additionally, we report FS and foot height of the human motions generated by DAViD, where we generate the same number of samples as in the test dataset. We demonstrate that DAViD generates more natural human motions compared to the baselines, with lower FS and foot height, as shown in Tab. 1.
Ablation Study. We conduct a quantitative ablation study on our key design choices, TGS and LoRA [31]. As shown in Tab. 1, we find that our contact guidance in TGS is crucial for expressing detailed hand contact in HOI motions. Note that this is not obvious, as contact guidance automatically detects potential contact region and guides them to get closer, and not provide information on where the contact should occur. The result shows that the contact guidance in TGS performs fine refinement of object motion within a learned distribution. We also perform an ablation study on our smoothness guidance in TGS by generating object motion using human motion from FullBodyManipulation [50] and measure the jitter. We define jitter following MocapEvery [46] and compute it for object translation. As shown in Tab. 2 (left), our smoothness guidance in TGS significantly reduces the object jitter that occurs when sampling object poses framwise, bringing it to a level similar to the jitter found in real-world capture data.
@@ 5 Discussion
In this paper, we present DAViD, a method to learn Dynamic Affordance, addressing that there exists not only static patterns (e.g., contact, spatial relation) but also dynamic patterns in HOIs. While obtaining a 4D HOI dataset including diverse objects is challenging, we observe that the video diffusion model has prior knowledge of the dynamic patterns for general objects. Compared to learning affordance from videos generated from text-to-video models, our pipeline, which first renders the 3D object to generate HOI images and then uses an image-to-video model, has the advantage of lifting affordances in video into 3D space. Using the 4D HOI samples generated through our lifting pipeline, we train DAViD, which consists of (1) the LoRA module of MDM for learning human motion patterns and (2) an object pose diffusion model with human pose condition. We qualitatively show that the 4D HOI samples generated by DAViD effectively model the dynamic pattern of human pose and corresponding object pose during HOIs. Quantitative results show that our DAViD models human motion and interactions that are more similar to real-world HOI than the baseline. In addition, we demonstrate the advantage of using LoRA to model the dynamic patterns of using specific objects by generating HOI motion that include multi-object interactions.
@@ Acknowledgements
This work was supported by NAVER Webtoon, NRF grant funded by the Korean government (MSIT) (No. RS-2022-NR070498, No. RS-2023-00218601), and IITP grant funded by the Korea government (MSIT) [No. RS-2024-00439854, RS-2021-II211343]. H. Joo is the corresponding author.
@@ B.1 Additional Qualitative Results
We showcase additional qualitative results in Fig. S.3. In Fig. S.3(a), we show the results of training DAViD on additional object categories. Fig. S.3 shows the results of generating various HOI motions using our trained LoRA [31]. Through the results of generating various HOI motions, we demonstrate that our LoRA [31] faithfully learns the dynamic patterns during HOI.
@@ B.2 Additional Quantitative Results
We conduct an additional quantitative evaluation on our method estimating object motion from the given video using PnP [19, 48] and 2D tracking [35, 34]. As shown in Tab. 3, our method outperforms the existing 6D pose estimation method, FoundationPose [101] in predicting 6D object poses in a single view at BEHAVE sequences via the commonly used ADD metric. As the FoundationPose [101] leverages depthmap from RGBD camera for 6D pose estimation, and our method only need RGB video, we use esimated depth for FoundationPose, which is from the metric depth model [4]. We also report the score of the best view for better assessment, as BEHAVE sequences contain object occlusion in a single view, which hinders estimating 6D object pose.
@@ B.3 Scale of the Object
[…]
Ours 43.40 64.96 85.52
$\text{FoundationPose}_{\text{w/ depth map estimation}}$ 24.67 29.91 40.89
Table 3: Object Pose Prediction against BEHAVE sequences.
@@ C.3 Modeling Dexterous Hand-Object Interaction
Although we model the human with SMPL-X [69], and recent 2D diffusion models demonstrate impressive quality in representing detailed hands, the pre-trained video diffusion model and 3D human estimator struggle to uplift the 2D hand from HOI images to high-quality 4D. This hinders the modeling of dexterous hand-object interactions in both our 4D HOI samples and the learned Dynamic Affordance. As a future direction, we can explore separately learning the hand patterns and merging them with the dynamic patterns we learned, with the expectation of improving the hand quality of the sampled HOI motion.

### v3

@@ 4.3 Qualitative Results
4D HOI Generation for Various Objects. Different from previous studies, we leverage the pre-trained video diffusion model, which allows to learn dynamic patterns of HOI for unbounded object categories. Fig. 4 visualizes the HOI motion for various objects generated by DAViD. The frames are visualized in temporal order. The results show that our method effectively models the dynamic patterns in HOIs. In Fig. 4 (a), Fig. 4 (b), Fig. 4 (c) we observe that the human motion and the relative position of the object are well modeled in relatively static scenarios. In cases with more dynamic human motion, we find that various interaction patterns, including pulling-Fig. 4 (d), pushing-Fig. 4 (h), lifting-Fig. 4 (g), and holding-Fig. 4 (f) are well aligned with the typical use of objects of humans. Also, we can observe that our method can effectively capture complex human motions and their corresponding object motions, such as back and forth movement, as shown in Fig. 4 (e) and Fig. 4 (j). Importantly, unlike previous studies where motion translation typically resulted from walking, we observe that our model effectively captures cases of translation changes without foot movement, as shown in Fig. 4 (i).
Figure 5: Combining Concepts. Our HOI motion concepts can be combined with original knowledge of pre-trained MDM, even for multiple HOI motion concepts.
@@
Methods $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1 Score}\uparrow$ $\text{FS}\downarrow$ $\text{H}_{\text{feet}}\downarrow$
@@
Methods $\text{FID}\downarrow$ $\text{Diversity}\rightarrow$
@@ 4.4 Quantitative Results
Comparison with Baselines. To verify the efficacy of our model, we compare DAViD against OMOMO [53] and CHOIS [54] on the FullBodyManipulation Dataset [53]. Specifically, we measure the performance of DAViD on each of the test datasets of the FullBodyManipulation [53], selected from baselines. For the metrics about HOI quality, hand contact is considered when the average distance between the 25 hand joints of SMPL-X [72] and the object vertices is smaller than a threshold, 0.05. As shown in Tab. 1, DAViD outperforms the baselines for all contact metrics. This shows that DAViD generates plausible HOI motion, demonstrating the efficacy of our approach in modeling the object pose distribution from the human pose, and sampling object motion using TGS. Additionally, we report FS and foot height of the human motions generated by DAViD, where we generate the same number of samples as in the test dataset. We demonstrate that DAViD generates more natural human motions compared to the baselines, with lower FS and foot height, as shown in Tab. 1.
Ablation Study. We conduct a quantitative ablation study on our key design choices, TGS and LoRA [33]. As shown in Tab. 1, we find that our contact guidance in TGS is crucial for expressing detailed hand contact in HOI motions. Note that this is not obvious, as contact guidance automatically detects potential contact region and guides them to get closer, and not provide information on where the contact should occur. The result shows that the contact guidance in TGS performs fine refinement of object motion within a learned distribution. We also perform an ablation study on our smoothness guidance in TGS by generating object motion using human motion from FullBodyManipulation [53] and measure the jitter. We define jitter following MocapEvery [49] and compute it for object translation. As shown in Tab. 2 (left), our smoothness guidance in TGS significantly reduces the object jitter that occurs when sampling object poses frame-wise, bringing it to a level similar to the jitter found in real-world capture data.
@@ 5 Discussion
In this paper, we present DAViD, a method to learn Dynamic Affordance, addressing both static patterns (e.g., contact, spatial relation) and dynamic motion patterns in HOIs. While obtaining a 4D HOI dataset including diverse objects is challenging, we observe that the video diffusion model has prior knowledge of the dynamic patterns for general objects. Compared to learning affordance from videos generated from text-to-video models, our pipeline, which first renders the 3D object to generate HOI images and then uses an image-to-video model, has the advantage of lifting affordances in video into 3D space. Using the 4D HOI samples generated through our lifting pipeline, we train DAViD, which consists of (1) the LoRA module of MDM for learning human motion patterns and (2) an object pose diffusion model with human pose condition. We qualitatively show that the 4D HOI samples generated by DAViD effectively model the dynamic pattern of human pose and corresponding object pose during HOIs. Quantitative results show that our DAViD models human motion and interactions that are more similar to real-world HOI than the baseline. In addition, we demonstrate the advantage of using LoRA to model the dynamic patterns of using specific objects by generating HOI motion that includes multi-object interactions.
@@ Acknowledgements
This work was supported by NAVER Webtoon, NRF grant funded by the Korean government (MSIT) (No. RS-2022-NR070498, No. RS-2023-00218601, and RS-2025-25396144), and IITP grant funded by the Korea government (MSIT) (No. RS-2024-00439854, RS-2021-II211343, and RS-2025-25442338). H. Joo is the corresponding author.
@@ B.1 Additional Qualitative Results
We showcase additional qualitative results in Fig. S.4 and Fig. S.5. In Fig. S.4, we show the results of generating various HOI motions using our trained DAViD. Through the results of generating various HOI motions, we demonstrate that our LoRA [33] faithfully learns the dynamic patterns during HOI. In Fig. S.5, we show the qualitative results generated from DAViD, trained on the FullBodyManip [53] dataset. We demonstrate that DAViD is not only able to learn coherent and simple HOI patterns, but also capable of generating relatively complex HOI motions.
Figure S.5: Qualitative Results on FullBodyManip Datatset. We showcase additional qualitative results of DAViD trained on the FullBodyManip dataset.
@@ B.2 Additional Quantitative Results
We report additional comparisons with our baselines for each category of the FullBodyManip dataset in Tab. 3. As our TGS automatically detects potential contact points and guides them closer during sampling, we vary the threshold of potential contact $\rho$ to examine the effect of our guidance. Note that the potential contact threshold $\rho$ used in TGS is lower than the threshold used for evaluating the metric (0.05), to ensure that points not considered as contact by the model are not forced into contact. As shown in Tab. 3, our contact guidance in TGS significantly improves recall and consequently the F1 score. We demonstrate that our contact guidance allows to sample fine-grained hand-object contact on the coarse distribution learned by our object pose diffusion model. In contrast, precision tends to remain stable or decrease as the threshold level increases, which appears to be a side effect of unintended potential contacts detected in the early stage of the denoising preocess. Empirically, we find that maintaining a low threshold around 0.02 minimizes side effects and is effective for sampling fine-grained hand-object contact.
Methods Clothesstand Floorlamp Largebox Largetable Monitor Plasticbox Smallbox
$\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$ $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$ $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$ $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$ $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$ $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$ $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$
$\text{DAViD}_{\rho=0.00}$ 0.667 0.088 0.156 0.611 0.297 0.400 0.894 0.578 0.702 0.917 0.500 0.647 0.807 0.38 0.517 0.934 0.463 0.619 0.988 0.464 0.632
$\text{DAViD}_{\rho=0.01}$ 0.913 0.309 0.462 0.500 0.486 0.493 0.862 0.311 0.457 0.743 0.197 0.311 0.847 0.485 0.617 0.848 0.275 0.415 0.985 0.515 0.676
$\text{DAViD}_{\rho=0.02}$ 0.765 0.382 0.510 0.531 0.459 0.493 0.889 0.497 0.637 0.929 0.598 0.728 0.799 0.442 0.570 0.865 0.553 0.675 0.986 0.508 0.671
$\text{DAViD}_{\rho=0.03}$ 0.536 0.221 0.313 0.600 0.243 0.346 0.863 0.547 0.669 0.837 0.545 0.661 0.823 0.523 0.640 0.821 0.471 0.599 0.982 0.545 0.701
CHOIS 0.615 0.353 0.449 0.667 0.378 0.483 0.773 0.211 0.332 0.783 0.136 0.232 0.827 0.156 0.263 0.674 0.119 0.202 0.957 0.239 0.382
$\text{DAViD}_{\rho=0.00}$ 0.667 0.098 0.171 0.784 0.138 0.235 0.951 0.485 0.642 0.808 0.371 0.509 0.849 0.345 0.491 0.881 0.349 0.500 0.991 0.528 0.689
$\text{DAViD}_{\rho=0.01}$ 0.429 0.197 0.270 0.615 0.114 0.193 0.979 0.539 0.695 0.839 0.173 0.287 0.895 0.488 0.632 0.824 0.371 0.512 0.990 0.610 0.755
$\text{DAViD}_{\rho=0.02}$ 0.405 0.279 0.330 0.667 0.276 0.391 0.954 0.488 0.645 0.873 0.456 0.599 0.911 0.602 0.725 0.800 0.446 0.572 0.989 0.649 0.784
$\text{DAViD}_{\rho=0.03}$ 0.571 0.328 0.417 0.472 0.119 0.190 0.959 0.588 0.729 0.868 0.533 0.661 0.877 0.564 0.687 0.790 0.522 0.629 0.990 0.645 0.781
OMOMO 0.432 0.311 0.362 0.828 0.114 0.201 0.917 0.617 0.738 0.868 0.603 0.711 0.776 0.432 0.555 0.831 0.342 0.484 0.982 0.639 0.774
Methods Smalltable Suitcase Trashcan Tripod Whitechair Woodchair Average
$\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$ $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$ $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$ $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$ $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$ $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$ $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{F1}\uparrow$
$\text{DAViD}_{\rho=0.00}$ 0.926 0.482 0.634 0.990 0.564 0.718 0.961 0.508 0.665 0.714 0.195 0.306 0.706 0.213 0.328 0.913 0.328 0.483 0.848 0.389 0.524
$\text{DAViD}_{\rho=0.01}$ 0.953 0.527 0.679 0.992 0.609 0.755 0.952 0.496 0.652 0.727 0.312 0.436 0.808 0.187 0.303 0.716 0.276 0.398 0.847 0.383 0.511
$\text{DAViD}_{\rho=0.02}$ 0.915 0.691 0.788 0.979 0.609 0.751 0.916 0.628 0.745 0.731 0.247 0.369 0.825 0.609 0.701 0.779 0.495 0.605 0.850 0.517 0.634
$\text{DAViD}_{\rho=0.03}$ 0.907 0.717 0.801 0.967 0.649 0.777 0.947 0.893 0.919 0.667 0.208 0.317 0.826 0.591 0.689 0.711 0.448 0.55 0.812 0.508 0.614
CHOIS 0.944 0.431 0.592 0.942 0.251 0.397 0.580 0.165 0.257 0.917 0.143 0.247 0.483 0.062 0.110 0.717 0.224 0.341 0.760 0.221 0.330
$\text{DAViD}_{\rho=0.00}$ 0.936 0.457 0.614 0.989 0.420 0.590 0.970 0.561 0.711 0.839 0.173 0.287 1.00 0.255 0.407 0.922 0.371 0.529 0.891 0.350 0.490
$\text{DAViD}_{\rho=0.01}$ 0.931 0.502 0.653 0.986 0.501 0.664 0.963 0.630 0.761 0.902 0.306 0.457 0.963 0.265 0.416 0.810 0.406 0.540 0.856 0.392 0.526
$\text{DAViD}_{\rho=0.02}$ 0.920 0.639 0.755 0.976 0.536 0.692 0.957 0.760 0.847 0.949 0.347 0.508 1.00 0.378 0.548 0.865 0.490 0.618 0.867 0.488 0.616
$\text{DAViD}_{\rho=0.03}$ 0.945 0.706 0.809 0.985 0.552 0.708 0.956 0.699 0.807 0.967 0.321 0.482 0.974 0.388 0.555 0.774 0.516 0.619 0.856 0.499 0.621
OMOMO 0.870 0.470 0.610 0.974 0.660 0.787 0.739 0.500 0.596 0.824 0.225 0.354 0.967 0.296 0.453 0.779 0.433 0.557 0.830 0.434 0.552
Table 3: Additional Quantitative Results. We report additional quantitative results for each category of the FullBodyManip dataset by varying the contact loss threshold used in our TGS.
@@ B.3 Scale of the Object
B.4 Generalizability Across Input 3D Objects.
By leveraging pre-trained 2D diffusion models, our 4D HOI sample generation pipeline is scalable to various object categories and instances. As shown in Fig. S.7, our pipeline allows to generate 4D HOI samples not only from 3D objects in existing datasets, but also from those reconstructed from in-the-wild images. For the given in-the-wild image, we first reconstruct 3D objects from the image via TRELLIS [110], using them as input to our pipeline to generate the 4D HOI sample interacting with the object.
@@ C.3 Modeling Dexterous Hand-Object Interaction
Although we model the human with SMPL-X [72], and recent 2D diffusion models demonstrate impressive quality in representing detailed hands, the pre-trained video diffusion model and 3D human estimator struggle to uplift the 2D hand from HOI images to high-quality 4D. This hinders the modeling of dexterous hand-object interactions in both our 4D HOI samples and the learned Dynamic Affordance. As a future direction, we can explore separately learning the hand patterns and merging them with the dynamic patterns we learned, with the expectation of improving the hand quality of the sampled HOI motion. As shown in Fig. S.6, we show that hand poses can be extended to our 4D HOI samples using a hand pose estimator [73] in a simple scenario.
@@ C.4 Concept Conflict
Figure S.6: Hand Pose Extension. DAVID uses SMPL-X as the human model, allowing hand pose extension in our 4D HOI samples.
Figure S.7: Generalizability Across Input 3D Objects. Our 4D HOI sample generation pipeline is generalizable to any input 3D object, including those reconstructed from images.